OpenStudy (lgbasallote):

Just checking an answer... Find the volume of the solid obtained by rotating the region bounded by \(y = \sqrt{25 - x^2}, \quad y = 0, \quad x = 2, \quad and \quad x = 4\) about the x - axis. is the answer \(74\pi\)??

OpenStudy (inkyvoyd):

\(\int_{2}^{4}\sqrt{25-x^2}dx\)?

OpenStudy (lgbasallote):

yep that's what i used...this problem is not washer right?

OpenStudy (anonymous):

i got something else... @lgbasallote

OpenStudy (lgbasallote):

aww

OpenStudy (inkyvoyd):

@dpaInc , did I do the reduction right?

OpenStudy (lgbasallote):

@inkyvoyd there was also a square

OpenStudy (inkyvoyd):

Yes, there was. :S

OpenStudy (anonymous):

isn't the integrand supposed to be f(x)^2 ?

OpenStudy (inkyvoyd):

\(\int_{2}^{4}-x^2+25 dx\)

OpenStudy (anonymous):

still, no... @inkyvoyd , that's it..

OpenStudy (anonymous):

don't forget the pi though...

OpenStudy (lgbasallote):

what did i do wrong =_=

OpenStudy (anonymous):

apple preferably....

OpenStudy (inkyvoyd):

So, \(\frac{-x^3}{3}+25x|_{x=2}^{x=4}\)

sam (.sam.):

I think you forgotten the pi too

OpenStudy (inkyvoyd):

No, I didn't. I made it invisible. -.-

OpenStudy (anonymous):

like himself... :)

OpenStudy (lgbasallote):

\[\pi | 25x - \frac{x^3}{3} |_2^4\] \[\pi {[(100 - 50) - (\frac{-64}{3} - \frac{8}{3})]}\]

OpenStudy (lgbasallote):

\[\pi [50 - (\frac{-72}{3})]\] \[\pi[50 + 24] = 74\pi\]

OpenStudy (lgbasallote):

I SEE WHAT I DID WRONG!!

OpenStudy (anonymous):

@inkyvoyd , why did you take out your soln?

OpenStudy (lgbasallote):

\[\frac{206\pi}{3}\] @dpaInc

OpenStudy (inkyvoyd):

Cause it was wrong.

OpenStudy (anonymous):

then mine is wrong too... 94pi/3 ???

OpenStudy (inkyvoyd):

I got like 32 and 1/3 pi

OpenStudy (inkyvoyd):

but that's wrong i think

OpenStudy (anonymous):

mine is 31 and 1/3 pi

OpenStudy (lgbasallote):

oh yeah lol 94pi/3 hahaha

OpenStudy (inkyvoyd):

@dpaInc is right "S

OpenStudy (anonymous):

change to spherical coordinates?

OpenStudy (anonymous):

you found your err?

OpenStudy (lgbasallote):

\[(100 - 50) - [\frac{64}{3} - \frac{8}{3}] = 50 - \frac{56}{3} = \frac{94}{3}\]

OpenStudy (anonymous):

oh, the 64 was -64 before....

OpenStudy (lgbasallote):

i double-negatived it haha

OpenStudy (anonymous):

i don't never do that... :)

OpenStudy (lgbasallote):

lol =)) well i gotta get home now..thanks :D

OpenStudy (lgbasallote):

(btw..im in the university and it's like 7 pm here getting creepy)

OpenStudy (anonymous):

roha...